LBTB

7:10
  • Seeing doesn't look very good. We'll have to see at around 8UT, if Jochen wants to do AO tonight.
7:51
  • Getting ready to switch over to AO
7:58
  • Several of the computers in the dome seem to be having thermal issues. The DIMM hasn't been working for a while and now we've lost RA and DEC information.
8:31
  • Presetting Bin3, 416 Hz GOPT might be good enough. 0.13
  • Gains 4.57 1.51 0.10, Mag 15.2
  • Going to run NCPA.
  • 20230702_083458
  • Seeing is somewhat variable, but the DIMM is defunct.
  • Glory be! It lives! 0.84, not too shabby.
8:42
  • Heading to a new target.
  • Bin 3, 988 Hz
  • Gains 5.48 0.72
  • GOPT is 0.21
  • Also going to use NCPA
  • 20230702_084900
9:04
  • Presetting to a Siefert.
  • Bin 3, 789 Hz. 14.8 Mag
  • GOPT ratio too small.
9:08
  • We have to unwrap the azimuth axis.
9:16
  • Reset the preset
  • 790 Hz Gains 4.21 2.30
  • 14.8 mag
  • 20230702_091810
  • GOPT ratio 0.02, no no no
9:56
  • New Preset
  • Bin3, 252 Hz
  • Gain g_tt_4.31_ho1_0.43_ho2_0.10.fits
  • TT is flat with these sieferts, meaning that I have to manually ramp it up due to the autogain wanting to set it to 0.16.
  • mag 16.5
  • 20230702_101722
10:45
  • We had some odd behavior from the camera lens and pupil check. It's as if the pupil check got confused and railed the camera lens, but in the wrong direction. some manual motion seems to have sorted it out.
  • I think it was due to the flux dropping as we have some patchy clouds right now.
11:19
  • Done with AO. It takes 65 years for the dumb OCAM to warm up.
-- GregoryTaylor - 02 Jul 2023
Topic revision: r1 - 02 Jul 2023, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback